OpenStudy (anonymous):

Help!!! Solving Elimination 4x+8y=20 4x+2y=-6

OpenStudy (anonymous):

i am so confused omgg what grade is this?

OpenStudy (texaschic101):

4x + 8y = 20 4x + 2y = -6 --->(-1) -------------- 4x + 8y = 20 -4x - 2y = 6 (result of multiplying by -1) --------------add 6y = 26 y = 26/6 which reduces to 13/3 now sub 13/3 in for y in either equation 4x + 8y = 20 4x + 8(13/3) = 20 4x + 104/3 = 20 4x = 20 - 104/3 4x = 60/3 - 104/3 4x = - 44/3 x = (-44/3) / 4 x = -44/3 * 1/4 x = -44/12 which reduces to -11/3 check.. 4x + 2y = -6 4(-11/3) + 2(13/3) = -6 -44/3 + 26/3 = -6 - 18/3 = -6 -6 = -6 (correct) solution is : (-11/3, 13/3)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

@lala 8 grade lol yeah I know it soo Hard :/ & yay a Thanks @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Lol omg I wrote it wrong -.- It's 4x+8y=20 4x+2y=-16 Sorry ;/